The controller of Standard Industries has collected the following monthly expense data for analyzing the cost behavior of electricity costs.

Total Total Electricity Costs Machine Hours
January $2,500 300
February   3,000 350
March   3,600 500
April   4,500 690
May   3,200 400
June   4,900 700
July   4,100 650
August   3,800 520
September   5,100 680
October   4,200 630
November   3,300 350
December   6,100 720

Instructions

(a)  

Determine the fixed- and variable-cost components using regression analysis (We recommend the use of Excel.)

(b)  

Prepare a scatter plot using Excel. Present the cost equation line estimated in part (a).

(c)  

What electricity cost does the cost equation estimate for a level of activity of 500 machine hours? By what amount does this differ from March's observed cost for 500 machine hours?
Answer #1

A)

x

y

  

Calculated y

  

Error

1.
2611.379321 111.3793209
2.
2904.864583 95.13541699
3.
3785.320369 185.3203693
4.
4900.564365 400.5643652
5.
3198.349845 1.650154894
6.
4959.261418 59.26141766
7.
4665.776156 565.7761556
8.
3902.714474 102.7144741
9.
4841.867313 258.1326872
10.
4548.382051 348.3820507
11.
2904.864583 395.135417
12.
5076.655522 1023.344478

Result:  y = 5.869705242 x + 850.4677484

Correlation Coefficient: r = 9.100540652·10-1
Residual Sum of Squares: rss = 1941787.566
Coefficient of Determination: R2 = 8.281984016·10-1

B) Booki - Excel (Product Activation Failed) ? * - - Sign in Doser FILE HOME INSERT PAGE LAYOUT * Cut Calibri - 11 Copy Paste Fo

C)

Here is the cost data that was observed:

Hours Cost
Highest Observation 720 $6100
Lowest Observation (300) (2500)
Difference 420 3600

Slope or Variable Cost: 3600/420 = $8.57

Y Intercept or Fixed Cost: $71

which makes the cost function as Y=8.57X + 71

so for the electricity cost for 500 hours is Y=8.57(500) + 71 = $4356.

where actually observed it is $3600. So its differ by 4356-3600 = $756.